# Experimental Evaluation of a Two-Way Intelligibility Protocol for Human-LLM Interaction

### v2

@@ Abstract
Our interest is in the design of software systems involving a human-expert interacting–using natural language–with a large language model (LLM) on data analysis tasks. For complex problems, it is possible that LLMs can harness human expertise and creativity to find solutions that were otherwise elusive. On one level, this interaction takes place through multiple turns of prompts from the human and responses from the LLM. Here we investigate a more structured approach based on an abstract protocol described in Baskar et al. (2024) for interaction between agents. The protocol is motivated by a notion of “two-way intelligibility” and is modelled by a pair of communicating finite-state machines. We provide an implementation of the protocol, and provide empirical evidence of using the implementation to mediate interactions between an LLM and a human-agent in two areas of scientific interest (radiology and drug design). We conduct controlled experiments with a human proxy (a database), and uncontrolled experiments with human subjects. The results provide evidence in support of the protocol’s capability of capturing one- and two-way intelligibility in human-LLM interaction; and for the utility of two-way intelligibility in the design of human-machine systems.
@@ 1 Introduction
For over four decades, researchers in human–machine systems have recognised the importance of making machine-generated predictions intelligible to humans. Michie Michie (1982) was among the first to note potential mismatches between human and machine representations. He proposed a classification of machine learning (ML) systems into three types Michie (1988): Weak ML systems are concerned only with improving performance, given sample data. Strong ML systems improve performance, but are also required to communicate what it has learned in some human-comprehensible form (Michie assumes this will be symbolic). Ultra-strong ML systems are Strong ML systems that can also teach the human to improve his or her performance.
This framework, intended for human-in-the-loop contexts, focuses on intelligibility, the ability of the system to explain the “what” and “why” within the human’s window of understanding rather than intelligence. The system’s internal model can use any representation, so long as communication is understandable. Michie’s scheme addresses one-way communication from machine to human, without metrics for intelligibility (later addressed in Ai et al. (2021)) or provisions for human-to-machine communication, which is important in collaborative discovery. This categorisation has also recently informed a similar 3-way categorisation for the use of AI tools in scientific discovery Krenn et al. (2022).
Despite growing interest in human–AI interaction Wienrich and Latoschik (2021); Gammelgård-Larsen et al. (2024); Tsiakas and Murray-Rust (2024); Guan (2024), few models prioritise mutual intelligibility. Explainable AI (XAI) often produces explanations without tailoring them to the intended audience, undermining user confidence when advice is opaque and when the system cannot gauge the user’s comprehension. A recent proposal Baskar et al. (2024) addresses this by defining “two-way intelligibility” as a property of communication protocols between agents, modelled as finite-state machines. Messages are tagged with one of four “R’s”: ratification, refutation, revision, or rejection, and intelligibility is assessed from the sequence of these tags. While Baskar et al. (2024) specifies the protocol with proofs of correctness and termination, no implementation or empirical evidence was provided.
Given the capabilities of large language models (LLMs) to produce human-readable predictions and explanations when properly conditioned, they are promising candidates for such protocols. However, determining the right conditioning information is non-trivial. This paper addresses that gap by implementing the Baskar et al. (2024) protocol restricted to interactions between an LLM-based generator-agent and a human tester-agent, providing a partial but practical realisation of two-way intelligibility in action.
@@ 2 Related Work
Iterative Prompting: Work has been done in prompting LLMs iteratively, by decomposing the problem into smaller sub-problems, or paths in a decision tree, such as Chain-of-Thought Wei et al. (2023); Sun et al. (2024), Tree-of-Thought Yao et al. (2023), etc. Our work is not a method to iteratively prompt an LLM; rather, it is a method to analyse and interpret the interactions between a Human and an LLM in a multi-turn fashion, using the notion of intelligibility.
Intelligibility: Intelligibility is the ability to be understood or comprehended. Previous work has proposed using concepts in philosophy, psychology, and cognitive science Miller (2018); Hoffman et al. (2022) to gain insight into intelligible systems. Our aim is to measure intelligibility and identify Strong/Ultra Strong ML systems as described by Michie Michie (1988), and see if the characteristics are shared across tasks. Our implementation allows us to explore interactions between not only Human-ML interactions but also interactions between two ML agents. We achieve this by restricting the Agents to Predict and Explain (PEX) as described in Baskar et al. (2024). This constraint enables us to quantify intelligibility using message tags.
@@ 3 The Main Aspects of the PXP Protocol
The PXP protocol is an interaction model described in Baskar et al. (2024). It is motivated by the question of when an interaction between agents could be said to be intelligible to either. Restricting attention to specific kinds of agents–called PEX agents–the paper is concerned with when predictions and explanations for a data instance produced by one PEX agent are intelligible to another PEX agent. The agents are modelled as finite-state machines, and “intelligibility” is defined as a property inferrable from the tagged messages exchanged between the finite-state machines.
An agent $a_{n}$ decides the tag of its message to be sent to agent $a_{m}$ based on comparing its prediction $y_{n}$ and explanation $e_{n}$ for a data instance $x$ to the prediction and explanation it received from $a_{m}$ for $x$ ($y_{m}$ and $e_{m}$ respectively). The comparisons define the truth-value of guards for transitions, which in turn define the message-tags. Informally, if $a_{n}$ agrees with $a_{m}$ on both prediction and explanation, it sends a message with a $\mathit{RATIFY}$} tag. If $a_{n}$ disagrees with either the prediction or explanation but not both, then it might change (“revise”) its (internal) model. If it can revise its model, then its message has a $\mathit{REVISE}$} tag; else it has a $\mathit{REFUTE}$} tag. Finally, if $a_{n}$ disagrees with both prediction and explanation, it sends a $\mathit{REJECT}$} tag. In all cases, the complete message has the tag, along with $a_{n}$’s current prediction and explanation. Intelligibility is then defined as properties inferable from observing the message-tags exchanged. The important definitions in Baskar et al. (2024) are:
@@
Two-way intelligibility follows directly if a session $S$ is one-way intelligible for $m$ and one-way intelligible for $n$. The authors in Baskar et al. (2024) then use these definitions to suggest a Michie-style categorisation of Strong and Ultra-Strong Intelligibility. Based on this suggestion, we define the following:
@@ 4 Modelling Human-LLM Interaction
The implementation described here is restricted to interaction between a single human-agent and a single machine-agent. The protocol in Baskar et al. (2024) is akin to a plain-old-telephone-system (POTS), and it is sufficient for our purposes to implement it as a rudimentary blackboard system with a simple scheduler that alternates between the two agents. The blackboard consists of 3 tables accessible to the agents. The tables are: (a) $Data$. This is a table consisting of $(s,x)$ pairs where $s$ is a session ID, and $x$ is a data instance; and (b) $Message$. A table of 5-tuples $(s,j,\alpha,\mu,\beta)$ where: $s$ is a session identifier, $j$ is a message-number, $\alpha$ is a sender-id, $\mu$ is a message and $\beta$ is a receiver-id; and (c) $Context$. This a table consisting of 3-tuples $(s,j,c)$ where $s$ is a session-id, $j$ is the message number, and $c$ is some domain-specific context information. For simplicity, message-numbers will be assumed to be from the set $\{0,1,2,\ldots\}$; $\alpha,\beta$ are from $\{h,m\}$ where $h$ denoting “human” and $m$ denotes “machine”; and messages $\mu$ are $(l,y,e)$ tuples, where $l$ is from $\{{\mathit{RATIFY}},{\mathit{REFUTE}},{\mathit{REVISE}},{\mathit{REJECT}}\}$, and $y$ and $e$ are the prediction and explanation respectively. We treat the blackboard as a relational database $\Delta$ consisting of the set of tables $\{D,M,C\}$. For presentability, $\Delta$ is denoted as a “shared” input in the agent-functions below.
@@ 5 Experimental Evaluation
5:while ($Left\neq\emptysetdo$
@@
In this section, we examine Human-LLM interaction through the use of the INTERACT procedure. Experiments reported here look at two real-world problems: X-Ray Diagnosis (RAD), and Molecule Synthesis (DRUG). In the former, we want to use LLMs for diagnosing X-rays and producing reports, and in the latter, we want proposals for synthesis pathways for molecules. We report on two kinds of experiments. First, controlled experiments are conducted using a database of human-authored predictions and explanations and an LLM. This allows us to perform repeated (simulated) experiments, which are needed since sampling variations can arise with the use of LLMs. Secondly, we report on results obtained in uncontrolled experiments using human subjects with varying levels of expertise interacting with an LLM. We use the experiments to assess the following: Human- and Machine-Intelligibility. We estimate the proportion of interactions exhibiting:(a) one- and two-way intelligibility; and(b) Strong- and Ultra-Strong Intelligibility. Machine-Performance. We estimate the changes in machine performance to increase when the interaction is at least one-way intelligible for the agent.111 Human-Performance cannot be assessed with controlled experiments, as the human proxy is a database that is not updated. Human-Performance can be assessed in uncontrolled experiments but it is difficult to establish baselines for comparative changes between specialists in complex domains with differing expertise. The results from uncontrolled experiments allow us to examine evidence to show that the simulation results are representative of real-life usage of the PXP protocol.
@@ 5.1 Problems
The RAD problem is concerned with obtaining predictions and explanations for up to 5 diseases from X-ray images. In this paper, an LLM is used to generate diagnoses and reports, given image data. We use data from the Radiopedia database Radiopaedia (2007). The DRUG problem is concerned with obtaining predictions and explanations for the synthesis of small molecules. Once potential molecules (leads) have been identified for inhibiting a target protein, we are interested in synthesising these molecules for testing on biological samples. In this paper, the DRUG problem examines the use of LLMs to propose plans for synthesis, and an explanation of why the plan is proposed. We use data from DrugBank Wishart et al. (2018). Dataset details are provided in Appendix B, and excerpts of sessions demonstrating the protocol’s use in Appendix D.
@@ 5.3 Method
The controlled experiments consist of 5 repetitions with 20 instances (X-ray images for RAD, and small molecules for DRUG). Uncontrolled experimental results are provided with DRUG (radiologists were unavailable). In Baskar et al. (2024), it is assumed that agents employ a LEARN function to decide whether to revise their predictions, based on estimating if revision would improve performance. In controlled experiments, we incorporate this through the use of a validation-set. Thus, the LLM only revises its prediction if its validation-set accuracy improves. In contrast, the human-proxy functions as an oracle (predictions and explanations are always taken to be correct). In uncontrolled experiments, no such set is used by either agent. The INTERACT procedure requires a bound on the total messages exchanged, and is set to $10$. We will use the usual (maximum-likelihood) estimate of proportion as the ratio of the number of sessions with a property to the total number of sessions; We use Claude 3.5 Sonnet as the LLM, with more details in Appendix A.
@@ 5.4 Results
We focus first on the controlled experiments. For simplicity, we will continue to call the human-proxy agent as “the human agent”. The statistics of interaction on controlled experiments are tabulated in Tab. 1, Fig. 12. Broadly, they are consistent with the following claims: (a) The proportion of one-way and two-way intelligible sessions for the human agent increases as the length of interaction increases. (b) Machine-performance (measured on test data) increases with the length of interaction. The following additional observations have possibly more interesting long-term consequences of using the $\tt{INTERACT\,}$} implementation of PXP : One- and Two-Way Intelligibility: The frequencies of sessions that are 1-way intelligible are high for both human- and machine-agents in RAD, and very high in DRUG. Some of this is due simply to the vast store of prior data and information within an LLM that allows it to provide correct predictions and explanations almost immediately. For example, in 8 of 20 sessions in DRUG, the chemist and machine-agent agree on the prediction and explanations within 3 exchanges. However, interestingly, in an additional 7 sessions, they agree after some exchange of refutations and revisions. For human-in-the-loop systems, this is indicative of the advantage to the human of being provided with explanations in natural language. It is also evidence that the LLM is able to act on text-based feedback from the human, consistent with the findings in Brown et al. (2020). The number of sessions that are 2-way intelligible clearly cannot be more than the lower number of 1-way intelligible sessions. Strong and Ultra-Strong Intelligibility: While it is not surprising to find the number of strongly intelligible sessions is lower than the number of 1- or 2-way intelligible sessions, the difference in numbers between strong- and ultra-strong sessions is surprisingly high. Resolving this requires examining first the sessions exhibiting strong intelligibility. Closer examination shows that these are exactly those sessions which are immediately ratified by both human- and machine-agents [^2] For all of these the message-tag sequence is: $\langle{\mathit{INIT}}_{m},{\mathit{RATIFY}}_{h},{\mathit{RATIFY}}_{m}\rangle$.. This is a degenerate form of strong-intelligibility (see Defn. 2). More interesting, strong-intelligibility would arise from longer sessions (for example, having a hypothetical sequence of message-tags: $\langle{\mathit{INIT}}_{m},{\mathit{REVISE}}_{h},$ ${\mathit{REVISE}}_{m},{\mathit{RATIFY}}_{h},{\mathit{RATIFY}}_{m}\rangle$). In fact, no such sequences occur. Thus, if we ignore these very short interactions, there are, in fact, very few strongly intelligible sessions. It is interesting though that in DRUG, we are able to observe 18 such sessions for the machine-agent.Variability. The experiments were repeated 5 times. It is also evident from the error bars that sampling variation decreases as the length of sessions increases. We believe this to be due to the increasing context information being available to the generator LLM used by the machine-agent as session-length increases.
@@
We turn now to results from the uncontrolled experiments, shown in Fig. 2(b). Broadly, we see the main trend obtained in controlled experiments repeated here: The proportion of intelligible sessions. It is interesting however, to note that the human-agent with higher computational–not chemical–expertise has a greater proportion of intelligible sessions. Closer inspection showed this to be a chance effect of the LLM’s output (the same molecule can produce very different synthesis plans from the LLM). Thus, while the LLM’s output remains broadly intelligible, the level of intelligibility can vary from one run to another. The Machine-Performance plot also brings out an interesting aspect. In the controlled experiments, the LLM’s prediction was only revised if performance on a validation-set improved. This check is not done in the uncontrolled setting, thus violating one of the assumptions of the protocol in Baskar et al. (2024). The result is that if the assumption is violated, then we can end up in the paradoxical position of the machine finding the human response intelligible, but that not being reflected in improvements in performance.

### v4

# Multi-Turn Human–LLM Interaction Through the Lens of a Two-Way Intelligibility Protocol

@@ Abstract
Our interest is in the design of software systems involving a human-expert interacting–using natural language–with a large language model (LLM) on data analysis tasks. For complex problems, it is possible that LLMs can harness human expertise and creativity to find solutions that were otherwise elusive. On one level, this interaction takes place through multiple turns of prompts from the human and responses from the LLM. Here we investigate a more structured approach based on an abstract protocol described in [3] for interaction between agents. The protocol is motivated by a notion of “two-way intelligibility” and is modelled by a pair of communicating finite-state machines. We provide an implementation of the protocol, and provide empirical evidence of using the implementation to mediate interactions between an LLM and a human-agent in two areas of scientific interest (radiology and drug design). We conduct controlled experiments with a human proxy (a database), and uncontrolled experiments with human subjects. The results provide evidence in support of the protocol’s capability of capturing one- and two-way intelligibility in human-LLM interaction; and for the utility of two-way intelligibility in the design of human-machine systems. Our code is available at https://github.com/karannb/interact.
@@ 1 Introduction
For over four decades, researchers in human–machine systems have recognised the importance of making machine-generated predictions intelligible to humans. Michie [10] was among the first to note potential mismatches between human and machine representations. He proposed a classification of machine learning (ML) systems into three types [11]: Weak ML systems are concerned only with improving performance, given sample data. Strong ML systems improve performance, but are also required to communicate what it has learned in some human-comprehensible form (Michie assumes this will be symbolic). Ultra-strong ML systems are Strong ML systems that can also teach the human to improve his or her performance.
This framework, intended for human-in-the-loop contexts, focuses on intelligibility, the ability of the system to explain the “what” and “why” within the human’s window of understanding rather than intelligence. The system’s internal model can use any representation, so long as communication is understandable. Michie’s scheme addresses one-way communication from machine to human, without metrics for intelligibility (later addressed in [1]) or provisions for human-to-machine communication, which is important in collaborative discovery. This categorisation has also recently informed a similar 3-way categorisation for the use of AI tools in scientific discovery [9].
Despite growing interest in human–AI interaction [17, 6, 15, 7], few models prioritise mutual intelligibility. Explainable AI (XAI) often produces explanations without tailoring them to the intended audience, undermining user confidence when advice is opaque and when the system cannot gauge the user’s comprehension. A recent proposal [3] addresses this by defining “two-way intelligibility” as a property of communication protocols between agents, modelled as finite-state machines. Messages are tagged with one of four “R’s”: ratification, refutation, revision, or rejection, and intelligibility is assessed from the sequence of these tags. While [3] specifies the protocol with proofs of correctness and termination, no implementation or empirical evidence was provided.
Given the capabilities of large language models (LLMs) to produce human-readable predictions and explanations when properly conditioned, they are promising candidates for such protocols. However, determining the right conditioning information is non-trivial. This paper addresses that gap by implementing the [3] protocol restricted to interactions between an LLM-based generator-agent and a human tester-agent, providing a partial but practical realisation of two-way intelligibility in action.
@@ 2 Related Work
Iterative Prompting: Work has been done in prompting LLMs iteratively, by decomposing the problem into smaller sub-problems, or paths in a decision tree, such as Chain-of-Thought [16, 14], Tree-of-Thought [19], etc. Our work is not a method to iteratively prompt an LLM; rather, it is a method to analyse and interpret the interactions between a Human and an LLM in a multi-turn fashion, using the notion of intelligibility.
Intelligibility: Intelligibility is the ability to be understood or comprehended. Previous work has proposed using concepts in philosophy, psychology, and cognitive science [12, 8] to gain insight into intelligible systems. Our aim is to measure intelligibility and identify Strong/Ultra Strong ML systems as described by Michie [11], and see if the characteristics are shared across tasks. Our implementation allows us to explore interactions between not only Human-ML interactions but also interactions between two ML agents. We achieve this by restricting the Agents to Predict and Explain (PEX) as described in [3]. This constraint enables us to quantify intelligibility using message tags.
@@ 3 The Main Aspects of the PXP Protocol
The PXP protocol is an interaction model described in [3]. It is motivated by the question of when an interaction between agents could be said to be intelligible to either. Restricting attention to specific kinds of agents–called PEX agents–the paper is concerned with when predictions and explanations for a data instance produced by one PEX agent are intelligible to another PEX agent. The agents are modelled as finite-state machines, and “intelligibility” is defined as a property inferrable from the tagged messages exchanged between the finite-state machines.
An agent $a_{n}$ decides the tag of its message to be sent to agent $a_{m}$ based on comparing its prediction $y_{n}$ and explanation $e_{n}$ for a data instance $x$ to the prediction and explanation it received from $a_{m}$ for $x$ ($y_{m}$ and $e_{m}$ respectively). The comparisons define the truth-value of guards for transitions, which in turn define the message-tags. Informally, if $a_{n}$ agrees with $a_{m}$ on both prediction and explanation, it sends a message with a $\mathit{RATIFY}$} tag. If $a_{n}$ disagrees with either the prediction or explanation but not both, then it might change (“revise”) its (internal) model. If it can revise its model, then its message has a $\mathit{REVISE}$} tag; else it has a $\mathit{REFUTE}$} tag. Finally, if $a_{n}$ disagrees with both prediction and explanation, it sends a $\mathit{REJECT}$} tag. In all cases, the complete message has the tag, along with $a_{n}$’s current prediction and explanation. Intelligibility is then defined as properties inferable from observing the message-tags exchanged. The important definitions in [3] are:
@@
Two-way intelligibility follows directly if a session $S$ is one-way intelligible for $m$ and one-way intelligible for $n$. The authors in [3] then use these definitions to suggest a Michie-style categorisation of Strong and Ultra-Strong Intelligibility. Based on this suggestion, we define the following:
@@ 4 Modelling Human-LLM Interaction
The implementation described here is restricted to interaction between a single human-agent and a single machine-agent. The protocol in [3] is akin to a plain-old-telephone-system (POTS), and it is sufficient for our purposes to implement it as a rudimentary blackboard system with a simple scheduler that alternates between the two agents. The blackboard consists of 3 tables accessible to the agents. The tables are: (a) $Data$. This is a table consisting of $(s,x)$ pairs where $s$ is a session ID, and $x$ is a data instance; and (b) $Message$. A table of 5-tuples $(s,j,\alpha,\mu,\beta)$ where: $s$ is a session identifier, $j$ is a message-number, $\alpha$ is a sender-id, $\mu$ is a message and $\beta$ is a receiver-id; and (c) $Context$. This a table consisting of 3-tuples $(s,j,c)$ where $s$ is a session-id, $j$ is the message number, and $c$ is some domain-specific context information. For simplicity, message-numbers will be assumed to be from the set $\{0,1,2,\ldots\}$; $\alpha,\beta$ are from $\{h,m\}$ where $h$ denoting “human” and $m$ denotes “machine”; and messages $\mu$ are $(l,y,e)$ tuples, where $l$ is from $\{{\mathit{RATIFY}},{\mathit{REFUTE}},{\mathit{REVISE}},{\mathit{REJECT}}\}$, and $y$ and $e$ are the prediction and explanation respectively. We treat the blackboard as a relational database $\Delta$ consisting of the set of tables $\{D,M,C\}$. For presentability, $\Delta$ is denoted as a “shared” input in the agent-functions below.
@@ 5 Experimental Evaluation
5:while ($Left\neq\emptyset$) do
@@
In this section, we examine Human-LLM interaction through the use of the INTERACT procedure. Experiments reported here look at two real-world problems: X-Ray Diagnosis (RAD), and Molecule Synthesis (DRUG). In the former, we want to use LLMs for diagnosing X-rays and producing reports, and in the latter, we want proposals for synthesis pathways for molecules. We report on two kinds of experiments. First, controlled experiments are conducted using a database of human-authored predictions and explanations and an LLM. This allows us to perform repeated (simulated) experiments, which are needed since sampling variations can arise with the use of LLMs. Secondly, we report on results obtained in uncontrolled experiments using human subjects with varying levels of expertise interacting with an LLM. We use the experiments to assess the following: Human- and Machine-Intelligibility. We estimate the proportion of interactions exhibiting:(a) one- and two-way intelligibility; and(b) Strong- and Ultra-Strong Intelligibility. Machine-Performance. We estimate the changes in machine performance to increase when the interaction is at least one-way intelligible for the agent. [^1] Human-Performance cannot be assessed with controlled experiments, as the human proxy is a database that is not updated. Human-Performance can be assessed in uncontrolled experiments but it is difficult to establish baselines for comparative changes between specialists in complex domains with differing expertise. The results from uncontrolled experiments allow us to examine evidence to show that the simulation results are representative of real-life usage of the PXP protocol.
@@ 5.1 Problems
The RAD problem is concerned with obtaining predictions and explanations for up to 5 diseases from X-ray images. In this paper, an LLM is used to generate diagnoses and reports, given image data. We use data from the Radiopedia database [13]. The DRUG problem is concerned with obtaining predictions and explanations for the synthesis of small molecules. Once potential molecules (leads) have been identified for inhibiting a target protein, we are interested in synthesising these molecules for testing on biological samples. In this paper, the DRUG problem examines the use of LLMs to propose plans for synthesis, and an explanation of why the plan is proposed. We use data from DrugBank [18]. Dataset details are provided in Appendix B, and excerpts of sessions demonstrating the protocol’s use in Appendix D.
@@ 5.3 Method
The controlled experiments consist of 5 repetitions with 20 instances (X-ray images for RAD, and small molecules for DRUG). Uncontrolled experimental results are provided with DRUG (radiologists were unavailable). In [3], it is assumed that agents employ a LEARN function to decide whether to revise their predictions, based on estimating if revision would improve performance. In controlled experiments, we incorporate this through the use of a validation-set. Thus, the LLM only revises its prediction if its validation-set accuracy improves. In contrast, the human-proxy functions as an oracle (predictions and explanations are always taken to be correct). In uncontrolled experiments, no such set is used by either agent. The INTERACT procedure requires a bound on the total messages exchanged, and is set to $10$. We will use the usual (maximum-likelihood) estimate of proportion as the ratio of the number of sessions with a property to the total number of sessions; We use Claude 3.5 Sonnet as the LLM, with more details in Appendix A.
@@ 5.4 Results
We focus first on the controlled experiments. For simplicity, we will continue to call the human-proxy agent as “the human agent”. The statistics of interaction on controlled experiments are tabulated in Tab. 1, Fig. 12. Broadly, they are consistent with the following claims: (a) The proportion of one-way and two-way intelligible sessions for the human agent increases as the length of interaction increases. (b) Machine-performance (measured on test data) increases with the length of interaction. The following additional observations have possibly more interesting long-term consequences of using the $\tt{INTERACT\,}$} implementation of PXP : One- and Two-Way Intelligibility: The frequencies of sessions that are 1-way intelligible are high for both human- and machine-agents in RAD, and very high in DRUG. Some of this is due simply to the vast store of prior data and information within an LLM that allows it to provide correct predictions and explanations almost immediately. For example, in 8 of 20 sessions in DRUG, the chemist and machine-agent agree on the prediction and explanations within 3 exchanges. However, interestingly, in an additional 7 sessions, they agree after some exchange of refutations and revisions. For human-in-the-loop systems, this is indicative of the advantage to the human of being provided with explanations in natural language. It is also evidence that the LLM is able to act on text-based feedback from the human, consistent with the findings in [4]. The number of sessions that are 2-way intelligible clearly cannot be more than the lower number of 1-way intelligible sessions. Strong and Ultra-Strong Intelligibility: While it is not surprising to find the number of strongly intelligible sessions is lower than the number of 1- or 2-way intelligible sessions, the difference in numbers between strong- and ultra-strong sessions is surprisingly high. Resolving this requires examining first the sessions exhibiting strong intelligibility. Closer examination shows that these are exactly those sessions which are immediately ratified by both human- and machine-agents [^2] For all of these the message-tag sequence is: $\langle{\mathit{INIT}}_{m},{\mathit{RATIFY}}_{h},{\mathit{RATIFY}}_{m}\rangle$.. This is a degenerate form of strong-intelligibility (see Defn. 2). More interesting, strong-intelligibility would arise from longer sessions (for example, having a hypothetical sequence of message-tags: $\langle{\mathit{INIT}}_{m},{\mathit{REVISE}}_{h},$ ${\mathit{REVISE}}_{m},{\mathit{RATIFY}}_{h},{\mathit{RATIFY}}_{m}\rangle$). In fact, no such sequences occur. Thus, if we ignore these very short interactions, there are, in fact, very few strongly intelligible sessions. It is interesting though that in DRUG, we are able to observe 18 such sessions for the machine-agent.Variability. The experiments were repeated 5 times. It is also evident from the error bars that sampling variation decreases as the length of sessions increases. We believe this to be due to the increasing context information being available to the generator LLM used by the machine-agent as session-length increases.
@@
We turn now to results from the uncontrolled experiments, shown in Fig. 2(b). Broadly, we see the main trend obtained in controlled experiments repeated here: The proportion of intelligible sessions. It is interesting however, to note that the human-agent with higher computational–not chemical–expertise has a greater proportion of intelligible sessions. Closer inspection showed this to be a chance effect of the LLM’s output (the same molecule can produce very different synthesis plans from the LLM). Thus, while the LLM’s output remains broadly intelligible, the level of intelligibility can vary from one run to another. The Machine-Performance plot also brings out an interesting aspect. In the controlled experiments, the LLM’s prediction was only revised if performance on a validation-set improved. This check is not done in the uncontrolled setting, thus violating one of the assumptions of the protocol in [3]. The result is that if the assumption is violated, then we can end up in the paradoxical position of the machine finding the human response intelligible, but that not being reflected in improvements in performance.